Determine the nature of the solutions of the equation. 2t^2-7t=0. Are there 2 imaginary solutions, 2 real solutions, or 1 real solution?
Answer by mananth(16946)

2t^2-7t=0
t(2t-7)=0
either t=0
OR
2t-7 =0
2t=7
t=7/2
two real solutions.